OpenStudy (anonymous):

Displacement vector d1 is in the yz plane 58.9 degrees from the positive direction of the y axis, has a positive z component, and has a magnitude of 4.52 m. Displacement vector d2 is in the xz plane 30.3 degrees from the positive direction of the x axis, has a positive z component, and has magnitude 1.63 m. What are (a) vector d1 dot product vector d2, (b) the x component of vector d1 and vector d2, the y component of vector d1 x vector d2, (d) the z component of vector d1 x vector d2, and (e) the angle between vector d1 and vector d2?

OpenStudy (anonymous):

@Michele_Laino could you help with this question? :)

OpenStudy (anonymous):

*me

OpenStudy (michele_laino):

Hint: here are your vectors, rewritten using the component notation: \[\begin{gathered} {{\mathbf{d}}_{\mathbf{1}}} = \left[ {0,\quad 4.52 \times \cos \left( {58.9} \right),\quad 4.52 \times \cos \left( {90 - 58.9} \right)} \right] \hfill \\ {{\mathbf{d}}_{\mathbf{2}}} = \left[ {1.63 \times \cos \left( {30.3} \right),\quad 0,\quad 1.63 \times \cos \left( {90 - 30.3} \right)} \right] \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):



OpenStudy (anonymous):

thank you :)